OpenStudy (anonymous):

What is the slope of the line that passes through the pair of points (3, –4), and (5, –7)? 2/3 -2/3 -3/2 3/2 I couldn't understand how to work this question out. @LogicalApple

OpenStudy (anonymous):

Set it up the same way as before. (x1, y1) = (3, -4) (x2, y2) = (5, -7) m = (y2 - y1) / (x2 - x1)

OpenStudy (anonymous):

But this time the only answers are fractions, and when i did it i got -1.5

OpenStudy (anonymous):

-1.5 is right :) But what is -1.5 as a fraction?

OpenStudy (anonymous):

-2/3?

OpenStudy (anonymous):

No wait, 3/2

OpenStudy (anonymous):

-3/2* -1 1/2

OpenStudy (anonymous):

-1.5 = -1 and 1/2. I know you are close

OpenStudy (anonymous):

You have to use the slope equation